OpenStudy (love_to_love_you):

A right cylindrical container perfectly holds 5 tennis balls placed one above the other. The radius of each ball is 3.35 centimeters. What is the approximate volume of the unoccupied space in the container? Use pi = 22 / 7

OpenStudy (noseboy908):

Hello there! To start, let's get the volume of each tennis ball. Do you know the volume formula for a sphere?

OpenStudy (love_to_love_you):

The volume is about 157.48

OpenStudy (love_to_love_you):

The formula is v = 4/3 pi r ^3

OpenStudy (noseboy908):

Correct, so do you have the volume of the cylinder as well?

OpenStudy (love_to_love_you):

That I do not

OpenStudy (love_to_love_you):

Would the radius be 3.35 for the cylinder too?

OpenStudy (noseboy908):

Correct. Your problem makes that part easy.

OpenStudy (love_to_love_you):

the height.. would I multiply the radius of the tennis balls by 5?

OpenStudy (noseboy908):

The volume for a cylinder is as follows:\[V = \pi r^2 h\]

OpenStudy (love_to_love_you):

So we know the radius, am I right about the height?

OpenStudy (love_to_love_you):

Or do we need to figure that out

OpenStudy (noseboy908):

We need the diameter, not the radius. That's an easy calculation, though.

OpenStudy (love_to_love_you):

6.7?

OpenStudy (noseboy908):

For each ball, yes. So multiply that by 5 to get the height.

OpenStudy (love_to_love_you):

33.5

OpenStudy (noseboy908):

Correct, now using that as height, input that with your radius into the volume formula above.

OpenStudy (love_to_love_you):

I got 1181.09?

OpenStudy (noseboy908):

Don't doubt yourself, you're correct. Note that that's the total volume of the cylinder containing the tennis balls. So let's take the total volume of each ball, as determined above, multiply that by 5, and subtract that from the the 1181.09 to get the empty space!

OpenStudy (love_to_love_you):

About 394 cm^3

OpenStudy (noseboy908):

Correct!

OpenStudy (love_to_love_you):

Thank you! Could you help me with another?

OpenStudy (noseboy908):

Sure.

OpenStudy (love_to_love_you):

Find the approximate volume of the cylinder shown in the diagram. Use pi = 22/7, and round your answer to the nearest integer value.

OpenStudy (love_to_love_you):

I suck at these, lol sorry

OpenStudy (noseboy908):

No worries, just give me a moment.

OpenStudy (love_to_love_you):

ok :)

OpenStudy (noseboy908):

Alright! Got it! For your own use, you may find this site very useful: http://goo.gl/A6Il02

OpenStudy (noseboy908):

Now, specifically, the site lists the formula for an oblique cylinder as \[V = hA(b)\] with the formula for A(b) being \[A(b) = \pi R^2\] Got it?

OpenStudy (love_to_love_you):

Yeah, okay

OpenStudy (noseboy908):

I know they're nothing simple, but taking it in steps makes it simple. What variables within the two formulas can we fill in right now?

OpenStudy (love_to_love_you):

the radius and height?

OpenStudy (noseboy908):

Yep. What are those?

OpenStudy (love_to_love_you):

R= 11 H= uhh 19?

OpenStudy (noseboy908):

Actually, I misspoke. We have to find the height, but that's a mere formula away! Do you recall the Pythagorean theorem?

OpenStudy (love_to_love_you):

Yes

OpenStudy (noseboy908):

The hypotenuse is easily known here, it's simply 19 (the slanted side). And we can gain the bottom value (what could be known as 'a') by taking the known radius, 11, and subtracting it from the bottom length of 19. What's that number?

OpenStudy (love_to_love_you):

8. But for the hypotenuse, do you mean 17 or 19?

OpenStudy (noseboy908):

The hypotenuse is 17cm, as given. Understand so far?

OpenStudy (love_to_love_you):

Yes. The height should be 15 then, right?

OpenStudy (noseboy908):

Exactly! (I really do appreciate you knowing how to solve formulas, you make my job really easy.) Now knowing h (15) and r (11), try to solve the question using the formulas above.

OpenStudy (love_to_love_you):

A(b)= 380.13 Is that correct?

OpenStudy (noseboy908):

Yep. Now multiply that by the height.

OpenStudy (love_to_love_you):

5701.95

OpenStudy (noseboy908):

Correct! And that's your final answer.

OpenStudy (love_to_love_you):

Awesome. Thank you again. You're a lifesaver! I might need help with another in a second if that's not too much trouble! I'll be sure to open a new question though so you can get the medal you deserve

OpenStudy (noseboy908):

Sure. Just mention me if I don't join immediately.

OpenStudy (love_to_love_you):

Alrighty :)